OpenStudy (anonymous):

4h+7h-16=6 4h+7h-16=6 @Mathematics

OpenStudy (anonymous):

can someone explain to me how they did this?

OpenStudy (anonymous):

11h=6+16 11h=22 h=22/11=2

OpenStudy (anonymous):

i dont understand what you did! ugh

OpenStudy (anonymous):

h = 2

OpenStudy (anonymous):

4h + 7h-16 = 6 11h = 6 + 16 11 h = 22 h = 22/11 h = 2

OpenStudy (katrinakaif):

Do you still need help with these?

OpenStudy (anonymous):

no its fine just that my teahcer is making us do it a diffrent way and thats what im confused on